OpenStudy (anonymous):

64a^2-16a^4=0 Solve equation

OpenStudy (anonymous):

-2, +2

OpenStudy (anonymous):

\[a=-2,+2\]

OpenStudy (anonymous):

\[16a^2(4-a^2)=0\] So solutions are: \[x=0, x=2, x=-1\]

OpenStudy (anonymous):

a^2=64/16 a^2=4

OpenStudy (anonymous):

i mean x= 0 , x= 2, x= -2

OpenStudy (anonymous):

oops, replace those x's with a's